Shift report 20011121 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Nov 21 2001 - 00:01:33 EST

  • Next message: Apache: "Shift report 20011121 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Nov 21 00:01:32 2001, Shift: 16:00-24:00
    Supervisor: J.I.Jordre, with: 
    Log book#11 pages 18 - 18
    
    Shift Summary:
    ------------------------------------------------------------
    1. Store from last shift ended at 19:45. Some cryo power supply had to 
       be replaced before they came up with new fresh beam at 22:00. 
       Still running.
    
    2. Run 5908 (from last shift) 	209k trigger 6
       Run 5910			127k trigger 6
       Run 5914 (during steering)	 84k trigger 6
       Run 5915 (still in process)	177k trigger 6
    
       Pedestal, DVM, TPC pedestal runs 5911-5913
    
    3. No new detector problems.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Nov 21 2001 - 00:02:28 EST